OpenStudy (noodya123):

someone please check my work and tell me if im wrong or correct Check all solutions to the system of inequalities X+y<=5+2 Y>2 Point A(2,5) Point B(1,1) Point C(5,2) Point D(3,6) Point E(2,-5) Point F(1,-1) point A 2+5<=5+2--->7<=7 y>2---->5>2 solution point B 1+1<=5+2----> 2<=7 y>2 ---->1>2 no solution point C 5+2<=5+2 --->7<=7 y>2---->2>2 no solution. Point D 3+6<=5+2---->9<=7 y>2----->6>2 no solution point E 2+(-5)<=5+2----> -3<=7 y>2----> -5>2 no solution 1+(-1)<=5+2-----> 0<=7 y>2----->-1>2 no solution so the only solution is

OpenStudy (radar):

I think that is only one problem......??? But I am not for sure lol

OpenStudy (radar):

There are multiple of points to verify.

OpenStudy (radar):

Would like to check your work, but I don't see where it is.

OpenStudy (radar):

I don't agree, with B. \[x + y \le 5 + 2\]\[x + y \le 7\]\[x \le 7 - y\]Checking Point B \[1 \le 7 - 1 \le 6\]Thats seems to be a solution. Do you agree?

OpenStudy (radar):

I see one more that I don't agree with.

OpenStudy (radar):

Also, I don't agree with your solution of the given problem. You have Y >=2 as the solution where I come up with \[Y \le 7 - X\]

OpenStudy (radar):

Good luck with your studies. I think you closed the question too early.